# Contrastive Alignment with Semantic Gap-Aware Corrections in Text-Video Retrieval

### v4

@@ D.2 Why Expanding with $\Delta_{ij}^{*}=0$ Is Inadequate
$\mathcal{L}_{i}(\Delta_{ij}^{*})\approx\mathcal{L}_{i}(0)+\sum_{j=1}^{B}\left[\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)\right]^{\top}\Delta_{ij},$ (17)
@@
D.3 The Need for a Non-Zero Initial $\Delta_{ij}^{*}$
@@
Thus, we need to initialize $\Delta_{ij}^{*}$ in a way that reflects the current state of all other pairs rather than assuming they are zero. The update for each $\Delta_{ij}$ should respect the relationship between all the pairwise similarities, as the optimization of one pair affects the relative similarity between all other pairs.
@@
D.4 Optimization Objective and Trust-Region Constrained Solution
@@
Having established the need for a coupled, non-zero initialization of $\Delta_{ij}$ to preserve the relative structure required by InfoNCE, we now formulate the optimization problem for finding the optimal update direction under a trust region constraint. Specifically, we assume a trust region radius $\varepsilon$ that bounds the magnitude of each $\Delta_{ij}$ and analyze the solutions in both the single-step (non-iterative) and iterative update settings.
Non-Iterative First-Order Update
@@
We consider the first-order Taylor expansion of the per-anchor loss $\mathcal{L}_{i}$ with respect to $\Delta_{ij}$ evaluated at the origin:
$\mathcal{L}_{i}(\Delta_{ij})\approx\mathcal{L}_{i}(0)+\nabla_{\Delta_{ij}}\mathcal{L}_{i}(0)^{\top}\cdot\Delta_{ij},$ (19)
subject to the trust-region constraint $\|\Delta_{ij}\|\leq\varepsilon$. This forms a standard constrained linear minimization problem, where the optimal $\Delta_{ij}$ minimizes the inner product with the gradient. By the Cauchy–Schwarz inequality:
$\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}^{\top}\cdot\Delta_{ij}\geq-\|\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}\|\cdot\|\Delta_{ij}\|,$ (20)
equality holds when $\Delta_{ij}$ and $-\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ are colinear. Thus, the optimal update is in the steepest descent direction:
$\Delta_{ij}^{*}=-\varepsilon\cdot\frac{\nabla_{\Delta_{ij}}\mathcal{L}_{i}}{\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}\|}.$ (21)
This solution, however, ignores the coupling between different $\Delta_{ik}$ and breaks the softmax structure, violating the relative comparison essential to InfoNCE, as discussed in the subsection above. Therefore, a more faithful solution must incorporate the influence of all $\Delta_{ij}$ jointly.
Iterative Update with Coupled Expansion
We instead consider the Taylor expansion centered at the current state $\Delta_{ij}^{(t)}$, capturing the coupled contributions of all other $\Delta_{ik}\neq\Delta_{ij}$. We seek the next update step:
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}+\delta,\quad\text{with }\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon.$ (22)
We again minimize the first-order approximation:
$\mathcal{L}_{i}(\Delta_{ij}^{(t+1)})\approx\mathcal{L}_{i}(\Delta_{ij}^{(t)})+\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}^{(t)}{}^{\top}\cdot(\Delta_{ij}^{(t+1)}-\Delta_{ij}^{(t)}).$
Letting $\hat{g}_{ij}^{(t)}=\nabla_{\Delta_{ij}}\mathcal{L}_{i}^{(t)}/\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}^{(t)}\|$, we propose a step in the steepest descent direction:
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\cdot\hat{g}_{ij}^{(t)}.$ (23)
To satisfy the norm constraint $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$, we solve for the maximum allowable step size $\alpha_{ij}^{(t)}$ using the triangle and Cauchy–Schwarz inequalities:
$\|\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\hat{g}_{ij}^{(t)}\|^{2}\leq\varepsilon^{2}.$ (24)
Expanding this yields the quadratic inequality:
$\alpha_{ij}^{(t)2}-2\alpha_{ij}^{(t)}\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\|\Delta_{ij}^{(t)}\|^{2}-\varepsilon^{2}\leq 0,$ (25)
whose maximal feasible solution is:
$\alpha_{ij}^{(t)}=\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\sqrt{\left(\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}\right)^{2}-\|\Delta_{ij}^{(t)}\|^{2}+\varepsilon^{2}}.$ (26)
Thus, the update maintains the trust region while moving in the steepest descent direction under the coupled loss landscape.
D.5 Why Use Neural Networks for Approximation?
The reason we opt for neural networks to approximate the current state of $\Delta_{ij}$ rather than directly solving for each $\Delta_{ij}$ through numerical optimization is primarily due to efficiency and scalability:
Handling Non-Linearities: The relationship between the embeddings of text and video is highly non-linear. Neural networks are particularly well-suited for learning such complex, non-linear mappings, allowing the model to effectively learn the best updates for $\Delta_{ij}$ without requiring explicit mathematical formulations for each pairwise interaction.
Scalability: As the size of the dataset increases, the computational cost of directly solving each $\Delta_{ij}$ grows rapidly. Neural networks allow for scalable learning since once trained, they can quickly compute updates for large batches without re-solving the optimization problem for each individual pair, making them a more practical choice in real-world settings.
In short, neural networks allow for faster convergence and better scalability compared to direct numerical solutions, making them essential for large-scale contrastive learning tasks.
Appendix E Computational Complexity Analysis.
In our proposed GARE framework, each mini-batch of size $B$ involves the prediction of a pair-specific increment $\Delta_{ij}\in\mathbb{R}^{D}$ for all text-video pairs. This yields a forward computational complexity of $\mathcal{O}(B^{2}D)$, since the learned increment network $\psi$ generates $B\times B$ outputs of dimension $D$. While quadratic in batch size, this cost is efficiently amortized during training and can be parallelized on modern hardware.
Importantly, this design avoids the substantial overhead required by a numerical optimization scheme. If one were to directly minimize the first-order Taylor expansion of the loss with respect to $\Delta$ using explicit iterative updates, two key obstacles arise: (1) The expansion is strictly first-order and lacks nonlinear modeling capacity, making it poorly suited for generalizing to unseen test-time structures; (2) Due to the softmax-coupled nature of InfoNCE, the gradients of each $\Delta_{ij}$ are entangled with all other $\Delta_{ik}$, requiring $T$ rounds of joint optimization for a converged update. As a result, the overall complexity escalates to $\mathcal{O}(TB^{2}D)$, where $T$ denotes the number of inner optimization steps. In addition to the increased cost, this setup suffers from severe gradient coupling and lacks a principled way to impose per-pair trust region constraints such as $|\Delta_{ij}|\leq\varepsilon$.
By contrast, our network-based approach amortizes this optimization and learns to implicitly enforce structure-aware constraints via training, making it not only more efficient but also more generalizable.

### v5

@@ H.2 Why Expanding with $\Delta_{i*}=0$ Is Inadequate
$\mathcal{L}_{i}(\Delta_{ij})\approx\mathcal{L}_{i}(0)+\nabla_{\Delta_{ij}}\mathcal{L}_{i}(0)^{\top}\Delta_{ij},$ (29)
@@
H.3 The Need for a Non-Zero Initial $\Delta_{i*}$
@@
Thus, we need to initialize $\Delta_{i*}$ in a way that reflects the current state of all other pairs rather than assuming they are zero. The update for each $\Delta_{ij}$ should respect the relationship between all the pairwise similarities, as the optimization of one pair affects the relative similarity between all other pairs.
@@
H.4 Optimization Objective and Trust-Region Constrained Solution
@@
Having established the need for a coupled, non-zero initialization of $\Delta_{ij}$ to preserve the relative structure required by InfoNCE, we now formulate the optimization problem for finding the approximate update direction under a trust region constraint. Specifically, we assume a trust region radius $\varepsilon$ that bounds the magnitude of each $\Delta_{ij}$ and analyze the solutions in both the single-step (non-iterative) and iterative update settings.
Non-iterative first-order update.
@@
We begin by considering the first-order Taylor expansion of the per-anchor loss $\mathcal{L}_{i}$ with respect to a single variable $\Delta_{ij}$ evaluated at the origin:
$\mathcal{L}_{i}(\Delta_{ij})\approx\mathcal{L}_{i}(0)+\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)^{\top}\Delta_{ij}.$ (31)
Since the constant term $\mathcal{L}_{i}(0)$ does not affect optimization, minimizing the approximated loss is equivalent to minimizing the linear term under a trust-region constraint:
$\min_{\|\Delta_{ij}\|\leq\varepsilon}\;\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)^{\top}\Delta_{ij},$ (32)
where $\varepsilon$ denotes the radius limiting the update magnitude of $\Delta_{ij}$. By the Cauchy–Schwarz inequality, for $g_{ij}=\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(0)$ we have
$g_{ij}^{\top}\Delta_{ij}\geq-\,\|g_{ij}\|\,\|\Delta_{ij}\|\geq-\,\|g_{ij}\|\,\varepsilon,$ (33)
which provides a tight lower bound for the objective within the trust region. The bound is attainable when $\Delta_{ij}$ is colinear with $g_{ij}$ and satisfies $\|\Delta_{ij}\|=\varepsilon$, yielding the feasible update
$\Delta_{ij}^{*}=-\,\varepsilon\cdot\frac{g_{ij}}{\|g_{ij}\|}.$ (34)
This closed-form solution achieves the maximal decrease in the first-order approximation of $\mathcal{L}_{i}$, corresponding to the steepest descent direction constrained by the trust region. However, this derivation treats $\mathcal{L}_{i}$ as a function of a single $\Delta_{ij}$ and neglects the coupling among $\{\Delta_{ik}\}_{k\neq j}$, thus breaking the relative similarity structure essential to InfoNCE. A more faithful formulation must therefore incorporate all $\Delta_{ij}$ jointly, as discussed in the following subsection.
Iterative update with coupled expansion.
We next analyze the update rule when $\Delta_{ij}$ is expanded around a coupled, non-zero state $\Delta_{i*}^{(t)}=\{\Delta_{ij}^{(t)}\}_{j=1}^{B}$, where each gradient $\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}$ is computed under the influence of all other nonzero $\Delta_{ik}^{(t)}$. We seek the next update step within a trust region:
$\Delta_{i*}^{(t+1)}=\Delta_{i*}^{(t)}+\delta,\quad\|\Delta_{i*}^{(t+1)}\|\leq\varepsilon.$ (35)
The first-order Taylor expansion of the per-anchor loss $\mathcal{L}_{i}$ at this coupled state is:
$\mathcal{L}_{i}(\Delta_{i*}^{(t+1)})\approx\mathcal{L}_{i}(\Delta_{i*}^{(t)})+\sum_{j=1}^{B}\big[\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})\big]^{\!\top}\big(\Delta_{ij}^{(t+1)}-\Delta_{ij}^{(t)}\big).$ (36)
This multivariate linearization serves to reveal that each partial gradient $\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})$ is conditioned on the current coupled state, and thus inherently encodes the interactions among all pairs. We do not minimize the linearized approximation itself; rather, it clarifies that the correct descent direction for each $\Delta_{ij}$ should be taken from the true gradient of $\mathcal{L}_{i}$ evaluated at the coupled state.
Following the principle of steepest descent, we update each component by
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\hat{g}_{ij}^{(t)},\qquad\hat{g}_{ij}^{(t)}=\frac{\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})}{\|\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})\|}.$ (37)
To satisfy the trust-region constraint $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$, we determine the maximal feasible step size $\alpha_{ij}^{(t)}$ from
$\|\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\hat{g}_{ij}^{(t)}\|^{2}\leq\varepsilon^{2},$ (38)
which expands to the quadratic inequality:
$\alpha_{ij}^{(t)2}-2\alpha_{ij}^{(t)}\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\|\Delta_{ij}^{(t)}\|^{2}-\varepsilon^{2}\leq 0.$ (39)
The maximal feasible solution is given by:
$\alpha_{ij}^{(t)}=\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\sqrt{\big(\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}\big)^{2}-\|\Delta_{ij}^{(t)}\|^{2}+\varepsilon^{2}}.$ (40)
This update ensures that each $\Delta_{ij}$ moves along its steepest descent direction while keeping its magnitude within the trust-region radius $\varepsilon$. Importantly, because all gradients $\nabla_{\!\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})$ are computed under the coupled multivariate state, the optimization preserves the relative ranking structure among all candidate pairs required by the InfoNCE objective.
H.5 Error Analysis
The iterative update of $\Delta_{ij}$ is derived from a multivariate first-order Taylor expansion of $\mathcal{L}_{i}$ under a trust-region constraint. This formulation defines a descent function space that guarantees loss reduction under the coupled gradient field, rather than prescribing a unique update direction for each $\Delta_{ij}$. Once the trust-region radius $\varepsilon_{ij}$ is specified (assuming an $\ell_{2}$-ball constraint), the steepest-descent direction becomes unique, and the theoretical update moves $\Delta_{ij}$ onto the boundary where the Cauchy–Schwarz inequality reaches equality. However, in practice, we do not explicitly assign a radius $\varepsilon_{ij}$ to each pair; instead, the effective trust-region constraint emerges implicitly from the model’s prior, which we regularize through a norm-based penalty that controls the magnitude of $\Delta$. Specifically, the $\Delta_{ij}$ produced by the $\psi$ network during the forward pass corresponds to the current state $\Delta_{ij}^{(t)}$, and the optimizer update during backpropagation embeds the next-step state $\Delta_{ij}^{(t+1)}$ into the parameters of $\psi$. When the model performs the next forward pass, the generated $\Delta_{ij}$ naturally reflects this updated state, while the norm-based regularization acts as the trust-region constraint on the previous update step. Therefore, the theoretical and practical procedures share the same steepest-descent direction, but their update magnitudes may differ due to the implicit learning of the trust-region radius through model priors. Consequently, our error analysis focuses on the consistency of update magnitudes between the theoretically derived trust-region step and the actual updates produced by the training framework.
Analytic step length. At the equality boundary of the trust-region constraint, the maximal feasible step size $\alpha_{ij}^{(t)}$ satisfying $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$ is given by:
$\alpha_{ij}^{(t)}=\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}+\sqrt{\big(\Delta_{ij}^{(t)\top}\hat{g}_{ij}^{(t)}\big)^{2}-\|\Delta_{ij}^{(t)}\|^{2}+\varepsilon^{2}},$ (41)
which projects the update exactly to the boundary of the feasible region and enforces norm-bounded motion of $\Delta_{ij}$.
Learned step length. During training, the neural module $\psi$ is optimized end-to-end via AdamW, implicitly learning how to adjust $\Delta_{ij}$ through gradient descent. After one step, the actual update can be expressed as:
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}-\eta_{ij}^{(t)}\frac{\nabla_{\!\Theta^{(t)}}\mathcal{L}_{i}^{(t)}}{\|\nabla_{\!\Theta^{(t)}}\mathcal{L}_{i}^{(t)}\|},$ (42)
where $\eta_{ij}^{(t)}$ denotes the effective step size produced by the optimizer. Although $\Delta_{ij}$ is updated through the training framework rather than by explicitly solving the analytic trust-region problem, we apply a norm-based trust-region regularization $\mathcal{L}_{\varepsilon}$ that restricts $\|\Delta_{ij}\|$ within a trust-region radius. This guarantees that the update magnitudes remain consistent with the theoretical constraint, even when learned implicitly.
To measure the deviation between the theoretical and learned step magnitudes, we compute the scalar error:
$E_{ij}^{(t)}=\|\eta_{ij}^{(t)}-\alpha_{ij}^{(t)}\|.$ (43)
This metric quantifies how closely the effective update magnitude in training adheres to the analytic trust-region step size.
Empirical analysis. Figure 9(a) and Figure 9(b) visualize the batch-wise mean error for positive and negative pairs, respectively. For positive pairs $(i=j)$, the error $\|\eta_{ii}-\alpha_{ii}\|$ remains within a stable range of $[1.0,4.5]$ and shows a clear convergence trend. Given the 512-dimensional embedding space, this corresponds to a per-dimension deviation of roughly $4.5/512$, which is negligible. This indicates that the learned $\Delta_{ii}$ updates remain consistent with the analytic trust-region dynamics and are well regularized by the norm constraint.
In contrast, for negative pairs $(i\neq j)$, the error $|\eta_{ij}-\alpha_{ij}|$ shows a divergent trend, which is theoretically expected. Each negative $\Delta_{ij}$ is influenced by repulsive gradients with (5121)(512-1) degrees of freedom, while the positive pair has a single dominant alignment direction. Moreover, since our norm-based regularization enlarges the updates whose magnitudes exceed the batch-wise mean $\|\Delta_{i*}\|$, the dispersion of negative $\Delta_{ij}$ naturally increases. Such divergence, however, plays a constructive role: by expanding negative $\{\Delta_{ij}\},(j\neq i)$ into a larger representational space, the model enhances embedding uniformity, allowing positive $\Delta_{ii}$ to achieve more stable alignment under a broader geometric margin.
(a) Batch-wise mean error for positive pairs $(i=j)$, showing convergence and bounded deviation ($[1.0,4.5]$).
(b) Batch-wise mean error for negative pairs $(i\neq j)$, exhibiting divergence that promotes embedding uniformity.
Figure 9: Comparison between the analytic step length $\alpha_{ij}$ and the effective AdamW step length $\eta_{ij}$. Although $\Delta_{ij}$ is learned through backpropagation, the norm-based trust-region constraint ensures that its magnitude remains bounded and consistent with the analytic formulation. Positive pairs show convergence and bounded errors, while negative pairs exhibit divergence that aids global uniformity.
Table 10: Comparison of computational efficiency between CLIP4Clip and GARE.
Metric CLIP4Clip GARE
Training time 1h 30m 1h 34m
Inference time 7.6s 6.9s
[…]
Inference memory (reserved) 4136 MB 4216 MB
Training FLOPs (per batch) 39,167.58 GFLOPs 39,287.55 GFLOPs (+0.3%)
Inference FLOPs (per batch) 11,868.70 GFLOPs 11,905.05 GFLOPs (+0.3%)
Table 11: Module-wise GFLOPs breakdown during forward propagation.
Module GFLOPs (%) Main computation source
CLIP visual encoder 11,673.60 (98.05%) ViT image processing
Video temporal transformer 38.81 (0.33%) FFN + self-attention
$\psi$ (cross-attention) 36.35 (0.31%) linear projection
CLIP text encoder 156.29 (1.31%) text sequence processing
Total 11,905.05 (100%)
Summary. In summary, our analysis demonstrates that while $\Delta_{ij}$ is optimized implicitly through the training framework, the norm-based trust-region constraint effectively maintains the magnitude of $\Delta$ within the theoretical bound. The learned $\psi$ module reproduces the analytic trust-region behavior for positive pairs, while the controlled divergence of negative pairs improves embedding uniformity—both consistent with the objectives of contrastive learning.
Inference Procedure (Pseudocode)
# Precompute all text and video embeddings offline using CLIP encoders
for each text batch T:
  for each video batch V:
    # Compute pair-specific delta via $\psi$ and adjust text embeddings
    logits = model.get_similarity_logits(T, V)
    # Append cosine-similarity block to the global similarity matrix
    sim_matrix.append(logits)
    # Discard the delta tensor immediately (transient variable)
# Concatenate all similarity blocks for final retrieval
Figure 10: Inference pseudocode of GARE. $\Delta$ are computed on-the-fly and discarded immediately to ensure constant memory usage during large-scale retrieval.
Cross-Attention Pairwise Parallelization (Simplified Code)
def cross_attention(query, key, value):
  # query: (a, b, dim) for (t_i - v_j), a is text batch size, b is video batch size
  # key/value: (b, f, dim) for video frames
  Q = Q_proj(query)
  K = K_proj(key)
  V = V_proj(value)
  logits = matmul(K, Q)
  scores = softmax(logits / sqrt(dim), dim=frame_dim)
  out = matmul(scores, V)
  return out
Figure 11: Simplified implementation of $\psi$’s cross-attention operation. The computation is pairwise-parallelizable across all text–video pairs, ensuring linear scaling with batch size and high GPU utilization.
Appendix I Model Efficiency and Implementation Details
GARE remains fully compatible with the standard dual-tower CLIP architecture while introducing only a lightweight cross-modal adjustment module, $\psi$. Implemented as a single-layer cross-attention transformer without FFN expansion, $\psi$ adds merely 1.58M parameters to the 354M parameters of the CLIP encoders. Its computational cost is negligible—only 36.35 GFLOPs per batch compared to CLIP’s 11,868.70 GFLOPs.
During inference, $\psi$ operates after both text and video embeddings have been precomputed, applying transient, pair-specific deltas before cosine similarity is calculated. This design preserves the precomputability and scalability of the dual-tower framework: all embeddings can be cached offline, $\psi$ performs only on-the-fly adjustments, and no delta tensors are stored—only similarity scores are retained.
Efficiency comparisons on MSR-VTT (4×RTX 4090 GPUs, batch size 128) show that GARE introduces minimal overhead (Table 10). A module-wise breakdown (Table 11) further confirms that over 98% of total FLOPs come from the visual encoder, while $\psi$ contributes less than 0.4%, demonstrating computational parity with the original CLIP.
As illustrated in Figure 10, GARE performs inference in a batch-parallel streaming manner, generating and discarding deltas on-the-fly to maintain constant memory usage. The internal cross-attention process (Figure 11) is pairwise-parallelizable, enabling simultaneous computation across all text–video pairs and ensuring high GPU utilization.
Overall, GARE retains the efficiency of dual-tower architectures while introducing a transient, pair-specific adjustment that enhances fine-grained cross-modal alignment without increasing latency or memory cost.
Distinction from traditional dual-Tower models.
Traditional dual-tower architectures define a consistent metric space where the triangle inequality holds, e.g., for any text–video pair $(t_{i},v_{j})$ and another video $v_{k}$,
$d(t_{i},v_{j})\leq d(t_{i},v_{k})+d(v_{k},v_{j}).$
This property allows efficient large-scale retrieval via approximate nearest neighbor (ANN) search, since all embeddings share a unified metric geometry.
In contrast, GARE introduces pair-specific adjustments through $\Delta_{ij}$, yielding a conditional distance
$d_{\text{GARE}}(t_{i},v_{j})=d(t_{i}+\Delta_{ij},,v_{j}),$
which depends on the paired video $v_{j}$. Substituting $t_{i}+\Delta_{ij}$ and $t_{i}+\Delta_{ik}$ into the above inequality breaks the shared metric assumption, and thus the triangle inequality no longer strictly holds. Consequently, while GARE improves fine-grained alignment, it cannot directly support ANN-based retrieval relying on fixed metric consistency.
From an efficiency standpoint, GARE computes $\Delta$ in a pair-wise parallel manner within each batch, discarding them immediately after computing cosine similarities. This design keeps latency low but introduces additional FLOPs that make direct large-scale application challenging. Nevertheless, we find that in practice the retrieved candidates from the unadjusted text embeddings $t_{i}$ already cover most of the top-ranked results produced by GARE. For instance, on the MSR-VTT 1k-A validation set, the top-256 candidates retrieved using $t_{i}$ include all of GARE’s top-10 matches.
Therefore, GARE naturally supports a two-stage retrieval pipeline: (1) use the dual-tower model with $t_{i}$ for large-scale ANN retrieval to obtain a compact candidate set, (2) re-rank the top-$k$ candidates using GARE’s pair-specific adjustments. This strategy preserves the scalability of dual-tower models while benefiting from GARE’s fine-grained alignment refinement.